If 16.0 g of Mg (OH)2 and 11.0 g of HCI are combined...what is the excess reactant?


1
Expert's answer
2021-02-28T06:21:18-0500

Mg(OH)2 + 2HCl = MgCl2 + 2H2O

For the reaction with 1 mol Mg(OH)2 0.5 mol HCl is needed

n = "\frac{m}{M}"

n(Mg(OH)2) = "\frac{16.0}{58,4} = 0.274 (mol)"

n(HCl) = "\frac{11}{36.5} = 0.31 (mol)"

"\frac{0.31}{2} = 0.165"

Answer: Mg(OH)2 is an excess reactant.